# Multimodal atmospheric super-resolution with deep generative models

### v1

@@ 2.5 Diffusion Data Assimilation for EDM
    Predict clean data: $\hat{\mathbf{x}}_{0}=F_{\theta}(\mathbf{x}_{\sigma_{i}};\sigma_{i})$.
@@
    Calculate prior drift: $\mathbf{d}_{\text{prior}}=(\mathbf{x}_{\sigma_{i}}-\hat{\mathbf{x}}_{0})/\sigma_{i}$.
@@
    Evaluate likelihood gradient: $\mathbf{s}_{l}(\mathbf{x}_{\sigma_{i}},\sigma_{i};\mathbf{y})$ using Eq. 13.
@@
    Calculate posterior drift: $\mathbf{d}_{\text{pos}}=\mathbf{d}_{\text{prior}}-\lambda_{g}\cdot\sigma_{i}\cdot\mathbf{s}_{l}$, where $\lambda_{g}$ is a guidance scale and $\sigma_{i}$ is multiplied to correctly incorporate it into the drift term of the probability flow ODE.
@@
    Perform sampling step: $\mathbf{x}_{\sigma_{i+1}}=\mathbf{x}_{\sigma_{i}}+\mathbf{d}_{\text{pos}}\cdot(\sigma_{i+1}-\sigma_{i})$.
@@
Algorithm 1 Stochastic Posterior Sampler
@@
2:Require: Churn schedule parameters $S_{\text{churn}},S_{t_{\min}},S_{t_{\max}}$.
@@
5:     $\gamma_{i}\leftarrow\min\left(\frac{S_{\text{churn}}}{N},\sqrt{2}-1\right)\text{ if }t_{i}\in[S_{t_{\min}},S_{t_{\max}}]\text{ else }0$triangleright Determine amount of noise injection (churn)
@@
8:     $\mathbf{x}_{\hat{t}_{i}}\leftarrow\mathbf{x}_{t_{i}}+\sqrt{\hat{t}_{i}^{2}-t_{i}^{2}}\boldsymbol{\epsilon}_{i}$triangleright Add noise to move from $t_{i}$ to $\hat{t}_{i}$
9:     $\hat{\mathbf{x}}_{0}^{\text{pred}}\leftarrow D_{\theta}(\mathbf{x}_{\hat{t}_{i}};\hat{t}_{i})$
10:     $\mathbf{d}_{\text{prior}}^{\text{pred}}\leftarrow(\mathbf{x}_{\hat{t}_{i}}-\hat{\mathbf{x}}_{0}^{\text{pred}})/\hat{t}_{i}$
11:     $\mathbf{s}_{l}^{\text{pred}}\leftarrow\nabla_{\mathbf{x}_{\hat{t}_{i}}}\log p(\mathbf{y}|\mathbf{x}_{\hat{t}_{i}})$triangleright Evaluate likelihood gradient using Equation 13
12:     $\mathbf{d}_{\text{pos}}^{\text{pred}}\leftarrow\mathbf{d}_{\text{prior}}^{\text{pred}}-\lambda_{g}\cdot\hat{t}_{i}\cdot\mathbf{s}_{l}^{\text{pred}}$
13:     $\mathbf{x}_{t_{i+1}}^{\text{pred}}\leftarrow\mathbf{x}_{\hat{t}_{i}}+\mathbf{d}_{\text{pos}}^{\text{pred}}\cdot(t_{i+1}-\hat{t}_{i})$triangleright Initial Euler step
14:     if $t_{i+1}\neq 0$ then
15:         $\hat{\mathbf{x}}_{0}^{\text{corr}}\leftarrow D_{\theta}(\mathbf{x}_{t_{i+1}}^{\text{pred}};t_{i+1})$
16:         $\mathbf{d}_{\text{prior}}^{\text{corr}}\leftarrow(\mathbf{x}_{t_{i+1}}^{\text{pred}}-\hat{\mathbf{x}}_{0}^{\text{corr}})/t_{i+1}$
17:         $\mathbf{s}_{l}^{\text{corr}}\leftarrow\nabla_{\mathbf{x}_{t_{i+1}}^{\text{pred}}}\log p(\mathbf{y}|\mathbf{x}_{t_{i+1}}^{\text{pred}})$triangleright Re-evaluate likelihood gradient using Equation 13 (optional)
18:         $\mathbf{d}_{\text{pos}}^{\text{corr}}\leftarrow\mathbf{d}_{\text{prior}}^{\text{corr}}-\lambda_{g}\cdot t_{i+1}\cdot\mathbf{s}_{l}^{\text{corr}}$
19:         $\mathbf{x}_{t_{i+1}}\leftarrow\mathbf{x}_{\hat{t}_{i}}+\frac{1}{2}(\mathbf{d}_{\text{pos}}^{\text{pred}}+\mathbf{d}_{\text{pos}}^{\text{corr}})\cdot(t_{i+1}-\hat{t}_{i})$triangleright 2nd-order correction step
20:     else
21:         $\mathbf{x}_{t_{i+1}}\leftarrow\mathbf{x}_{t_{i+1}}^{\text{pred}}$
22:     end if
23:end for
24:return $\mathbf{x}_{t_{N}}$
@@ 3.2 Data
We use ERA5, a state-of-the-art global reanalysis dataset produced by the European Centre for Medium-Range Weather Forecasts (ECMWF), which provides hourly estimates of a large number of atmospheric, land, and oceanic climate variables at a resolution of  31 km [23]. For this work, we use the 1.40525° (128 × 256 grid points) ERA5 dataset from WeatherBench [78], which used bilinear interpolation to regrid from the native resolution to 1.4°. The choice to use a lower resolution version of ERA5 was due to the computational limitations of training an unconditional diffusion model without large scale compute access.
For observational data, we rely on the Integrated Global Radiosonde Archive (IGRA), a comprehensive dataset of weather balloon observations from several stations worldwide, providing upper-air measurements like temperature, humidity, geopotential, pressure, and wind at different pressure levels [79]. Furthermore, we also construct an observation data set that is a further coarsened ERA5 at 11.242°(16x32 grid points) using bicubic interpolation. This is mentioned as LR further in the results.
@@ 3.3 Measurement Operators
For our super-resolution tasks, the measurement operation involves a downsampling operator that reduces the resolution of the data using interpolation techniques such as bicubic filtering. These downsampled values are compared with the available low-resolution data and used to guide our diffusion model. We mention these samples as Sample-LR. In contrast, our data assimilation tasks involve irregular or sparse observational locations. Our IGRA operator interpolates values at arbitrary spatiotemporal coordinates (e.g., latitude and longitude) using grid-based sampling. It defines a non-uniform measurement operator $M$ that evaluates the high-resolution field at the queried coordinates via bicubic interpolation in normalized coordinate space. This enables the comparison between model predictions and sparse ground truth measurements, forming the basis of our diffusion data assimilation frameworks. We mention these samples as Sample-IGRA. For our multimodal data assimilation we compute the misfit for both low-resolution and the IGRA measurements and use a weighted combination of the misfits to guide our diffusion sampling. We mention these samples as Sample IGRA+LR henceforth. For instance, the operator $M$ in Equation 13 can be the IGRA operator, which takes the generated full field as input and returns the corresponding values at the locations where we have the true IGRA measurements $\mathbf{y}$.
(a) 2m Temperature (K)
(b) U-Wind (m/s)
(c) V-Wind (m/s)
(d) Specific humidity (kg/kg)
@@ 4 Results and discussion
We outline results from our various experiments in the following. The training of the score-based generative model required approximately 30 hours on 16 Nvidia A100 GPUs, while generating a sample conditionally or with observation of sparse information requires only 10 seconds on 1 A100 GPU.
@@ 4.1 Super-resolution from structured coarse-grained fields
Figure 2: Ensemble based mean (left) and standard deviations (right) obtained from ensembles obtained with sample-LR zero-shot sampling. Showing 2m temperature, 10m u component of wind, 10m v component of wind, and specific humidity at 1000mb pressure from top to bottom.
@@
(d) 1000mb specific humidity
Figure 3: Spectral recovery exhibited by zero-shot sampling from ERA5 generative model when using sample-LR for zero-shot sampling.
@@
In the first experiment, we assess the ability to reconstruct the base-resolution ERA5 state given low-resolution observations from the same dataset (sample-LR). We qualitatively assess the reconstruction ability via test-time score-matching using visualizations of the fully recovered flow-fields for some selected variables in Figure 1 where coarse-grained sparse observations as well as the ground truth flow-fields at a particular instant are available. Error contours are obtained from using the ensemble mean as the predicted full state recovery. The contours indicate the ability to recover fine-scaled features using the score-matching technique effectively. We also assess the mean and variance of the generated samples from the score-matching procedure in Figure 2 where we can see correlations between variables and regions of high or low uncertainty. In particular, for recovering the 2m-temperature variable, we observe a higher uncertainty over the continents in contrast with wind speeds at 10m heights. For recovering specific humidity at 1000mb, we observe no coherent regions of high uncertainty aside from the general region of the lower latitudes where high values of specific humidity are commonly observed.
Quantitative analyses are also performed by looking at the quality of spectral recovery from angle-averaged kinetic energy spectra for the various variables as shown in Figure 3. In these plots, it can be seen that the recovered spectra from the zero-shot score-matching procedure leads to good agreement with the ground truth spectra for different variables all the way through to the cut-off wavenumber. Finally, we also assess the spatial variability of the generated fields by looking at traces along select latitudes across longitude in Figure 4. Here, the ensemble mean and variance are also plotted showing certain regions with higher uncertainty. In general, a good recovery of the trends is observed, including the capture of sharp peaks and troughs accurately across different variables.
@@ 4.2 Super-resolution from unstructured data and different modality
(a) Geopotential 200
(b) Specific humidity 1000
(c) Temperature 150
Figure 7: Scatters for additional fields at non-IGRA locations using zero-shot sampling from sample-IGRA. The knowledge of IGRA sensor measurements improves the reconstruction at sensor locations significantly as seen above.
@@
In Figure 6, we plot scatters for three representative variables at IGRA locations, as generated by the score-matching procedure. We expect to see the values generate collapse on a 45 line indicating that IGRA observations have been assimilated well. We also show the ground truth ERA5 values at the specific day in these scatters, where significant deviations are observed for the wind speeds. This indicates that the state recovery procedure has prioritized the likelihood obtained by IGRA observations at this point. In the same figure, we assess the performance of the generative model (after score-matching) for recovering the ground truth ERA5 at non-sampled locations given IGRA observations. Here it is seen that the spread for the generated values at non-IGRA points is generally reduced (in particular for temperature and specific humidity), indicating the IGRA observations have helped constrain the generator to the specific state. The spread from the unconditional samples indicates that while the unconditional generative process is able to provide fields that are ERA5-like, they do not recover the specific flow-fields without score-matching.

### v2

@@ 2.5 Diffusion Data Fusion for EDM
    Predict clean data: $\hat{\mathbf{x}}_{0}=D_{\theta}(\mathbf{x}_{t_{i}};t_{i})$.
@@
    Calculate prior drift: $\mathbf{d}_{\text{prior}}=(\mathbf{x}_{t_{i}}-\hat{\mathbf{x}}_{0})/t_{i}$.
@@
    Evaluate likelihood gradient: $\mathbf{s}_{l}(\mathbf{x}_{t_{i}},t_{i};\mathbf{y})$ using Eq. 13.
@@
    Calculate posterior drift: $\mathbf{d}_{\text{pos}}=\mathbf{d}_{\text{prior}}-\lambda_{g}\cdot t_{i}\cdot\mathbf{s}_{l}$, where $\lambda_{g}$ is a guidance scale and $t_{i}$ is multiplied to scale the drift term of the probability flow ODE.
@@
    Perform sampling step: $\mathbf{x}_{t_{i+1}}=\mathbf{x}_{t_{i}}+\mathbf{d}_{\text{pos}}\cdot(t_{i+1}-t_{i})$.
@@
Algorithm 1 Stochastic Posterior Sampler for data fusion
@@
2:Require: Churn schedule parameters $S_{\text{churn}},S_{t_{\min}},S_{t_{\max}}$. (refer [41])
@@
5:  $\gamma_{i}\leftarrow\min\left(\frac{S_{\text{churn}}}{N},\sqrt{2}-1\right)\text{ if }t_{i}\in[S_{t_{\min}},S_{t_{\max}}]\text{ else }0$ $\triangleright$ Determine amount of noise injection (churn)
@@
8:  $\mathbf{x}_{\hat{t}_{i}}\leftarrow\mathbf{x}_{t_{i}}+\sqrt{\hat{t}_{i}^{2}-t_{i}^{2}}\boldsymbol{\epsilon}_{i}$ $\triangleright$ Add noise to move from $t_{i}$ to $\hat{t}_{i}$
9:  $\mathbf{d}_{\text{prior}}^{\text{pred}}\leftarrow(\mathbf{x}_{\hat{t}_{i}}-D_{\theta}(\mathbf{x}_{\hat{t}_{i}};\hat{t}_{i}))/\hat{t}_{i}$
10:  $\mathbf{s}_{l}^{\text{pred}}\leftarrow\nabla_{\mathbf{x}_{\hat{t}_{i}}}\log p(\mathbf{y}|\mathbf{x}_{\hat{t}_{i}})$ $\triangleright$ Evaluate likelihood gradient using Equation 13
11:  $\mathbf{d}_{\text{pos}}^{\text{pred}}\leftarrow\mathbf{d}_{\text{prior}}^{\text{pred}}-\lambda_{g}\cdot\hat{t}_{i}\cdot\mathbf{s}_{l}^{\text{pred}}$
12:  $\mathbf{x}_{t_{i+1}}^{\text{pred}}\leftarrow\mathbf{x}_{\hat{t}_{i}}+\mathbf{d}_{\text{pos}}^{\text{pred}}\cdot(t_{i+1}-\hat{t}_{i})$ $\triangleright$ Initial Euler step
13:  if $t_{i+1}\neq 0$ then
14:   $\mathbf{d}_{\text{prior}}^{\text{corr}}\leftarrow(\mathbf{x}_{t_{i+1}}^{\text{pred}}-D_{\theta}(\mathbf{x}_{t_{i+1}}^{\text{pred}};t_{i+1}))/t_{i+1}$
15:   $\mathbf{s}_{l}^{\text{corr}}\leftarrow\nabla_{\mathbf{x}_{t_{i+1}}^{\text{pred}}}\log p(\mathbf{y}|\mathbf{x}_{t_{i+1}}^{\text{pred}})$ $\triangleright$ Re-evaluate likelihood gradient using Equation 13 (optional)
16:   $\mathbf{d}_{\text{pos}}^{\text{corr}}\leftarrow\mathbf{d}_{\text{prior}}^{\text{corr}}-\lambda_{g}\cdot t_{i+1}\cdot\mathbf{s}_{l}^{\text{corr}}$
17:   $\mathbf{x}_{t_{i+1}}\leftarrow\mathbf{x}_{\hat{t}_{i}}+\frac{1}{2}(\mathbf{d}_{\text{pos}}^{\text{pred}}+\mathbf{d}_{\text{pos}}^{\text{corr}})\cdot(t_{i+1}-\hat{t}_{i})$ $\triangleright$ 2nd-order correction step
18:  else
19:   $\mathbf{x}_{t_{i+1}}\leftarrow\mathbf{x}_{t_{i+1}}^{\text{pred}}$
20:  end if
21:end for
22:return $\mathbf{x}_{t_{N}}$
@@ 3.2 Data
We use ERA5, a state-of-the-art global reanalysis dataset produced by the European Centre for Medium-Range Weather Forecasts (ECMWF), which provides hourly estimates of a large number of atmospheric, land, and oceanic climate variables at a resolution of $\approx$31 km [23]. For this work, we use the 1.40525-degree (128 × 256 grid points) ERA5 dataset from WeatherBench [78], for which we used bilinear interpolation to regrid from the native resolution. The choice to use a lower resolution version of ERA5 was due to the computational limitations of training an unconditional diffusion model without large scale compute access.
For observational data, we rely on the Integrated Global Radiosonde Archive (IGRA), a comprehensive dataset of weather balloon observations from several stations worldwide, providing upper-air measurements like temperature, humidity, geopotential, pressure, and wind at different pressure levels [79]. Furthermore, we also construct an observation data set that is a further coarsened ERA5 at 11.242°(16x32 grid points) using bicubic interpolation. This is mentioned as ‘LR’ further in the results.
Next, we also assume access to a dynamical core comprised of a machine learning emulator for the atmosphere that can be used for guiding posterior sampling. We employ the Lightweight Uncoupled Climate Emulator (LUCIE) [80], which was trained on ERA5 reanalysis covering the period 1980–2000 at 3.75(48x96 grid points). LUCIE is trained on the same set of variables used in the diffusion model, with the addition of top-of-atmosphere incident solar radiation as an external forcing variable and makes forecasts at 6 hour resolutions. We emphasize that this dynamical core is a climate model, and therefore represents a source that generates observations from a fundamentally different platform than weather models. This is exemplified, for example, by the extremely coarse grid on which variables are resolved.
@@ 3.3 Measurement Operators
For our super-resolution tasks, we assume access to data on a coarse or sparse and unstructured grid from which the full state must be reconstructed. A first experiment involves the artificial construction of such data, from the ground truth, using a measurement operation. This measurement operation involves a downsampling operator that reduces the resolution of the data using interpolation techniques such as bicubic interpolation. These downsampled values are compared with the available low-resolution data and used to guide our diffusion model. We mention these samples as ‘Sample-LR’. In contrast, real-world data fusion tasks may involve irregular or sparse observational locations. Consequently, our IGRA operator interpolates values at arbitrary spatiotemporal coordinates (e.g., latitude and longitude) using grid-based sampling. It defines a non-uniform measurement operator $M$ that evaluates the high-resolution field at the queried coordinates via bicubic interpolation in normalized coordinate space. This enables the comparison between model predictions and sparse ground truth measurements, forming the basis of our diffusion data assimilation frameworks. We mention these samples as ‘Sample-IGRA’. For a first multimodal data fusion experiment, where we are performing super-resolution given access to both LR and IGRA information, we compute the misfit for both low-resolution and the IGRA measurements and use a weighted combination of the misfits to guide our diffusion sampling. We mention these samples as ‘Sample IGRA+LR’ henceforth. For instance, the operator $M$ in Equation 13 can be the IGRA operator, which takes the generated full field as input and returns the corresponding values at the locations where we have the true IGRA measurements $\mathbf{y}$.
Next, we detail how observations from the LUCIE atmospheric emulator may be introduced as an additional likelihood to guide posterior sampling while also observing IGRA measurements. In particular, the combined score function can be obtained by adding the individual scores with a hyperparameter $\lambda$ to adjust the weights given to the IGRA score.
[…]
where the subscript $L$ stands for LUCIE and $I$ stands for IGRA in each term. This likelihood score is used in Algorithm 1. Optimal choices of the $\lambda$ parameter require an ablation study which is provided in Section 4.3, Table 2.
@@ 3.4 Network architecture and hyperparameters
$\Sigma_{y}$ $\hat{\Gamma}$ 2m Temp RMSE (Mean $\pm$ Std)
$\times 10^{-1} 1×1$$\times 10^{-2}$ 2.3511 $\pm$ 0.1246
$\times 10^{-1} 1×1$$\times 10^{-3}$ 1.7273 $\pm$ 0.0529
$1\times 10^{-1}$ $5\times 10^{-3}$ 1.6883 $\pm$ 0.0224
$5\times 10^{-1}$ $1\times 10^{-3}$ 2.3673 $\pm$ 0.0834
$5\times 10^{-1}$ $5\times 10^{-3}$ 2.3359 $\pm$ 0.1129
$5\times 10^{-2}$ $5\times 10^{-3}$ 54.4904 $\pm$ 7.0866
Table 1: RMSE of 2m temperature from ERA5 for different $\Sigma_{y}$ and $\hat{\Gamma}$ settings. Mean $\pm$ Std is reported. Lowest RMSE mean is shown in bold.
3.5 Comparisons to similar frameworks
Our proposed framework is conceptually similar to the Appa-Weather approach [70] which also uses EDM-based posterior sampling. However, there are a few key differences in the present study. First, we study the assimilation of IGRA observations, as well as AI-emulator forecasts, in the reanalysis reconstruction state representing a data and model fusion exercise that leverages multiple sources of data generation that leverage both static and dynamical information about the evolving atmospheric state. Secondly, the emphasis of this paper is to assess the relative importance of these data in both the spatial and temporal sense while performing reconstructions. Compared to DiffDA [61], our approach is unconditional; therefore, our diffusion model does not need to be re-trained when new data or dynamical states become available. We also acknowledge previous efforts for data assimilation using diffusion models using observation data [62]. However, the notable difference in our work is the ability to perform multimodal data fusion with a low-resolution dynamical core and IGRA observations to obtain temporally consistent trajectories. Our approach, therefore, is a novel combination of super-resolution and data assimilation across multiple sources and resolutions of atmospheric observations.
@@ 4 Results and discussion
We outline results from our various experiments in the following. The training of the score-based generative model required approximately 30 hours on 16 Nvidia A100 GPUs, while generating a sample conditionally or with observation of sparse information requires only 2 seconds per ensemble on 1 A100 GPU. The results are generated for the year 2020 which is unseen data for training. We use Root Mean Square Error (RMSE) as the metric for comparing the diffusion samples with ERA5 and also for ablation studies.
@@ 4.1 Super-resolution from structured coarse-grained fields
(a) 500mb Geopotential ($m^{2}/s^{2}$)
(b) 10m U-Wind (m/s)
(c) 10m V-Wind (m/s)
(d) 850mb Specific humidity (kg/kg)
Figure 2: Comparison of 2m temperature fields from low-resolution input (LR), super-resolution using zero-shot posterior sampling (SR), and bicubic super resolution. The bottom row shows the true field and the corresponding error maps for SR and bicubic methods relative to the true field.
Figure 3: Comparison of 500mb V-wind fields from low-resolution input (LR), super-resolution using zero-shot posterior sampling (SR), and bicubic interpolation. The bottom row shows the true field and the corresponding error maps for SR and bicubic methods relative to the true field.
Figure 4: Ensemble based mean (left) and standard deviations (right) obtained from ensembles obtained with sample-LR zero-shot sampling. Showing 2m temperature, 10m u component of wind, 10m v component of wind, and specific humidity at 850mb pressure from top to bottom. One can observe a fine pattern of low uncertainty on a uniform grid, which corresponds to locations for the coarsely sampled ERA5.
@@
(d) 850mb specific humidity
Figure 5: Spectral recovery exhibited by zero-shot sampling from ERA5 generative model when using sample-LR for zero-shot sampling. Note how the proposed approach recovers the right spectral trend as against that of bicubic interpolation
@@
In the first experiment, we assess the ability to reconstruct the base-resolution ERA5 state given low-resolution observations from the same dataset (sample-LR). We qualitatively assess the reconstruction ability via test-time score-matching using visualizations of the fully recovered flow-fields for some selected variables in Figure 1 where coarse-grained sparse observations as well as the ground truth flow-fields at a particular instant are available. Error contours are obtained from using the ensemble mean as the predicted full state recovery. The contours indicate the ability to recover fine-scaled features using the score-matching technique effectively. We also assess the mean and variance of the generated samples from the score-matching procedure in Figure 4 where we can see correlations between variables and regions of high or low uncertainty. In particular, for recovering the 2m-temperature variable, we observe a higher uncertainty over the continents in contrast with wind speeds at 10m heights. For recovering specific humidity at 850mb, we observe no coherent regions of high uncertainty aside from the general region of the lower latitudes where high values of specific humidity are commonly observed.
Quantitative analyses are also performed by looking at the quality of spectral recovery from angle-averaged kinetic energy spectra for the various variables as shown in Figure 5(d). In these plots, it can be seen that the recovered spectra from the zero-shot score-matching procedure leads to good agreement with the ground truth spectra for different variables all the way through to the cut-off wavenumber. Additionally, we also assess the spatial variability of the generated fields by looking at traces along select latitudes across longitude in Figure 6. These latitudes are selected to reflect large variations that may be challenging for sample generation. Here, the ensemble mean and variance are also plotted showing certain regions with higher uncertainty. In general, a good recovery of the trends is observed, including the capture of sharp peaks and troughs accurately across different variables.
Next, we compare our super-resolution results against bicubic interpolation in Figures 2 and 3. It is clearly observed that the diffusion samples are far superior, both in accuracy and spatial consistency. We observe the same for the spectrum (Figure 5(d)) and traces along latitudes (Figure 6). We remark that the purpose of this experiment is not to propose a novel super-resolution algorithm for the atmosphere, of which there are several examples [81, 82, 83]. Instead, we assess the promise of super-resolution when interpreted as a posterior sampling process when likelihoods are generated from coarse data and the prior is generated from fine-scaled training data. Indeed, we note that most deep learning super-resolution frameworks build a direct map between coarse and fine-grid data which may provide more accurate reconstructions of the fine scales at the cost of flexibility of the coarse-grid representation. We remind the reader that the posterior sampling framework of diffusion allows one to perform reconstructions with various modalities and resolutions of data without retraining the diffusion model. These are demonstrated in the rest of the article.
@@ 4.2 Super-resolution from unstructured data and different modality
(a) Geopotential 500
(b) Temperature 850
(c) Humidity 850
Figure 9: Scatters for additional fields at non-IGRA locations using zero-shot sampling from sample-IGRA. The knowledge of IGRA sensor measurements improves the reconstruction mostly at sensor locations as seen above.
@@
In Figure 8, we plot scatters for three representative variables at IGRA locations, as generated by the score-matching procedure. We expect to see the values generate collapse on a 45 line indicating that IGRA observations have been assimilated well. We also show the ground truth ERA5 values at the specific day in these scatters, where significant deviations are observed for the wind speeds. This indicates that the state recovery procedure has prioritized the likelihood obtained by IGRA observations at this point. In the same figure, we assess the performance of the generative model (after score-matching) for recovering the ground truth ERA5 at non-sampled locations given IGRA observations. Here it is seen that the spread for the generated values at non-IGRA points is generally reduced (in particular for temperature and specific humidity), indicating the IGRA observations have helped constrain the generator to the specific state. The spread from the unconditional samples indicates that while the unconditional generative process is able to provide fields that are ERA5-like, they do not recover the specific flow-fields without score-matching. Furthermore, we also assess the reconstruction quality of the generative super-resolution for additional fields in Figure 9 at points that are not collocated with IGRA sensing locations. We observe that for some variables, such as geopotential height at 500mb and temperature at 850mb, IGRA observations lead to much improved reconstructions of ERA5. This represents a potential approach for rapid assessment of the quality of observations for recovering specific variables in the atmosphere.